If the graph of the line x-2y=-1 in the xy-plane is reflected across the x-axis, which of the following is an equation of the resulting reflection?
Answer by stanbon(75887)

If the graph of the line x-2y=-1 in the xy-plane is reflected across the x-axis, which of the following is an equation of the resulting reflection?
------
Reflection across the x-axis causes all the y-values to be multiplied by -1.
Your equation:
x-2y=-1
2y = x+1
y = (1/2)x+(1/2)
---
Your Problem:
x-2(-y)=-1
x+2y = -1
2y = -1-x
y = (-1/2)x-(1/2)
---
Graph of the two lines:

============================
Cheers,
Stan H.